# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ A.3 Proof of Theorem 2
$\frac{1}{t^{1-p}}\log\left(\frac{16t^{2}}{\delta}\right)=\frac{1}{t^{1-p}}\log\left(\frac{16}{\delta}\right)+\frac{2}{1-p}\frac{\log(t^{1-p})}{t^{1-p}}\leq\log\left(\frac{16}{\delta}\right)+\frac{2}{e(1-p)}.$
@@
$\leq\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
@@
$\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\leq 8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}$
@@
Proof of Theorem 2.
This is an immediate consequence of Lemmas 10 and 11. ∎
Appendix B Additional Results
B.1 Asymptotic convergence
We show that, even when $\sigma_{t}$ is not chosen as $\varsigma(t+1)^{-p}$ as in Theorems 1 and 2, we can still get almost sure convergence of our iterates for both Algorithms 1 and 2 (though without rates), as long as the regularization sequence satisfies the conditions in Lemma 4 or Lemma 9 respectively.
Theorem 3.
Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by either Algorithm 1 under the same conditions as Lemma 4. Then with probability $1$, we have
Proof of Theorem 3.
Lemma 4 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we may conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Propositions 1 and 2, we deduce that for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have, for all $t\geq 1$,
$F(x_{t})-F_{\operatorname{opt}}$ $\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
$\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left(\frac{8dt^{2}}{\delta}\right)}$
where the second inequality holds due to 1. We claim that all terms on the right hand side converge to 00 as $t\to\infty$. This is obvious except for the last term.
Recall that almost surely $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$. Therefore we have $\limsup_{t\to\infty}(F_{\operatorname{opt}}-F(x_{t}))\leq 0$, hence $\lim_{t\to\infty}\max\{F_{\operatorname{opt}}-F(x_{t}),0\}=0$ almost surely. Applying the Stolz–Cesàro theorem and (33) leads to
almost surely. Then each term on the right hand side of the inequality converges to 00 as $t\to\infty$. It follows that with probability at least $1-\delta$,
$\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})$ $\leq 0.$
Thus, since this holds for any $\delta\in(0,1)$, it holds almost surely. In conclusion, we have shown that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ and $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, which implies that $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely. ∎
Theorem 4.
Let $\{x_{t}\}_{t\geq 0}$ be the iterates generated by Algorithm 2 under the same conditions as Lemma 9. Then with probability $1$, we have
Proof of Theorem 4.
Lemma 9 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Proposition 1, Lemma 8 and Proposition 3, with probability at least $1-\delta$, we have
@@ B.1 Asymptotic convergence
Similar to the proof of Theorem 3, all terms on the right-hand side converge to 00 as $t\to\infty$, implying that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$. Combined with $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, this yields $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely, concluding the proof. ∎
B.2 Nonconvex setting
Define the auxiliary functions
$\mathcal{F}(x):=\max_{v\in X_{\operatorname{opt}}}\nabla F(x)^{\top}(x-v)\quad\&\quad\mathcal{G}(x):=G(x)-G_{\operatorname{opt}}$
for all $x\in X$. Recall that if $x\in X_{\operatorname{opt}}$ then $\mathcal{F}(x)=0$ is the necessary and sufficient optimality condition in convex case. Thus, $\mathcal{F}$ serves as a generalization of the stationarity condition $\nabla F(x)=0$ in the unconstrained setting. We start by establishing a bound, akin to Lemma 2, that holds even if $F$ is nonconvex.
Lemma 12.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with $\{\alpha_{t},\sigma_{t}\}_{t\geq 0}$ given by $\alpha_{t}=(t+1)^{-\omega}$ and $\sigma_{t}=\varsigma(t+1)^{-p}$ for $t\geq k$, with $0<p\leq\omega$. If 1 holds with the exception that $F$ may be nonconvex, then for every $t\geq k$, we have
Proof of Lemma 12.
Using the definition of $\Phi_{t}=\sigma_{t}F+G$, for any $t\geq k$, we have
where the inequality follows directly from (9) and by minimizing over $w_{t}\in X$. We emphasize that (9) does not involve any convexity assumptions on $\Phi_{t}$ and relies solely on its smoothness. Furthermore, notice that
$\leq\min_{v\in X_{\operatorname{opt}}}~{}\left\{\sigma\nabla F(x)^{\top}(v-x)+G(x)-G(v)\right\}$
$=-\sigma\mathcal{F}(x)-\mathcal{G}(x),$
[…]
Using the definition of $\alpha_{t}$, we arrive at
Unwinding this recursion back to $i=k$, we obtain
Since $p\leq\omega$, we have $(i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1}\geq 0$, therefore
$\sum_{i=k}^{t}(i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))$
$=2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k}).$
This completes the proof. ∎
B.2.1 One-sample stochastic setting
In this section we establish the convergence of Algorithm 1 in the nonconvex setting.
Theorem 5.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsizes $\alpha_{t}=(t\!+\!1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t\!+\!1)^{-p}$ for any $t\!\geq\!0$, with $p\!=\!\frac{2}{7}$ and $\omega\!=\!\frac{6}{7}$. If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{7}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
Proof of Theorem 5.
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
Applying Lemma 12 with $k=0$, with probability at least $1-\delta$, we thus have
[…]
Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
which further implies that
$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left(\frac{4d(t+q)^{2}}{\delta}\right)}.$
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Eq. 26, we have
$+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
we arrive at
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}$ $\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$
$+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$. We want the right hand side of both bounds $\to 0$. In order to guarantee this, we choose $p,\omega$ to minimize the slowest rate in terms of $t$:
which is realized by setting $p=2/7$, $\omega=6/7$ as required. ∎
B.2.2 Finite-sum setting
In this section we establish the convergence of Algorithm 2 in the nonconvex setting. To that end, we introduce the sequence $\{\beta_{t}\}_{t\geq 0}$, defined as
$\beta_{t}=\begin{cases}\displaystyle\frac{\sigma_{0}(q+1)^{\omega}\log(q+1)}{q+1}\left(1-\frac{\log(q+1)}{q+1}\right)^{q-t}&0\leq t\leq q,\\ \sigma_{t}&t>q.\end{cases}$ (27)
Theorem 6.
@@ B.2.2 Finite-sum setting
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$. If 1 holds with the exception that $F$ may be nonconvex and the sequence $\{\beta_{t}\}_{t\geq 0}$ is defined as in (27), then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
@@
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right)\quad\&\quad\mathcal{G}(x_{t})\leq O\left(t^{-\frac{1}{4}}\log\left(\frac{dt^{2}}{\delta}\right)\right).$
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}\!=\!\alpha$ for any $0\leq t\leq q$ and $\alpha_{t}=(t+1)^{-\omega}$ with $\omega>\frac{1}{2}$ for any $t>q$. If 1 holds with the exception that $F$ may be nonconvex, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (28)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
@@
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (29)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
@@
Proof of Lemma 13.
Recall from the proof of Lemma 7 that, employing the concentration inequality in [41, Theorem 3.5] yields
@@
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$ yields
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
@@
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
@@
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t\leq q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, for any $10\leq t\leq q+$ 1, it holds that
@@
Proof of Lemma 14.
Applying the bound (25) and after a straightforward re-arrangement, under 4, for any $0\leq i\leq t$, we obtain
@@
Given $\delta\in(0,1)$, we apply Lemma 13 to deduce that with probability at least $1-\delta/(i+1)(i+2)$, we have
@@
We finally analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha$ for any $0\leq t<q$, $\alpha_{t}=(t+1)^{-w}$ with $\omega>\frac{1}{2}$ for any $t\geq q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}\frac{1}{(i+1)(i+2)}$, for any $t>q$, it holds that
@@
Given $\delta\in(0,1)$ and $i>q$, by Lemma 13 and the inequality Eq. 31, with probability at least $1-\delta/(i+1)(i+2)$, we have
@@
We now instantiate the regularization parameter and analyze Algorithm 2.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\alpha:=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$. If 1 holds with the exception that $F$ may be nonconvex and 4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<t-1}1/(i+1)(i+2)$, for any $t>q+1$, it holds that
@@
Proof of Lemma 15.
By Lemma 12, for any $t\geq q+1$, we have
@@
Applying Lemma 14, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, we have
@@
Combining these two bounds with Proposition 4 then yields
@@
We are now well-equipped to prove Theorem 6.
Proof of Theorem 6.
We first focus on the terms involving $\mathcal{F}(x_{i})$ in Lemma 15. Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
$\textstyle-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i})$
@@
where $\alpha=\log(q+1)/(q+1)$, and the second and third inequalities follow from the geometric series bound and the definition of $\sigma_{t}$. Using this bound together with Lemma 15, we arrive at
$\mathcal{G}(x_{t})$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$
@@
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Lemma 15 and the definition of the sequence $\{\beta_{t}\}_{t\geq 0}$, we obtain
@@
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$

### v2

@@ C.1 Proof of Theorem˜3
$\frac{1}{t^{1-p}}\log\left(\frac{16t^{2}}{\delta}\right)=\frac{1}{t^{1-p}}\log\left(16/\delta\right)+\frac{2}{1-p}\frac{\log(t^{1-p})}{t^{1-p}}\leq\log\left(16/\delta\right)+\frac{2}{e(1-p)}.$
@@
$\leq\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)$
@@
$\frac{\left(8\sqrt{2}t-4q\right)(\sigma_{0}L_{f}+L_{g})D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)t\sigma_{t}}\leq 8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}$
@@
Proof of Theorem˜3.
The first part is an immediate consequence of Lemmas˜12 and 13. Thus, we devote this proof for the second part by proving that a slightly more general result that as long as 1 to 2 and ˜4 hold, the asymptotic convergence remains.
Lemma˜11 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Proposition˜1, Lemma˜10 and Proposition˜3, with probability at least $1-\delta$, we have
@@
Similar to the proof of Theorem˜1, all terms on the right-hand side converge to 0 as $t\to\infty$, implying that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ which holds with probability at least $1-\delta$ for any $\delta\in(0,1)$. Thus, $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ almost surely. Combined with $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, this yields $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely, concluding the proof. ∎
C.2 Proof of Theorem˜4
Here, we restate Theorem˜4 with exact upper bounds on stationary gaps.
Theorem  4.
@@ C.2 Proof of Theorem˜4
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with stepsizes $\alpha_{t}=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$, and regularization parameters $\sigma_{t}=\varsigma(\max\{t,q+1\}+1)^{-p}$ with $p=\frac{1}{2}$ and $w=\frac{3}{4}$. If ˜1 holds with the exception that $F$ may be nonconvex and the sequence $\{\beta_{t}\}_{t\geq 0}$ is defined as in (4), then with probability at least $1-\delta$, for every $t\geq q$, it (jointly) holds that
@@
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{G(x_{0})-G_{\text{opt}}}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}\left(a+bt^{1-p}\right)^{-1},$
$\mathcal{G}(x_{t})$ $\leq\left(\frac{G(x_{0})-G_{\text{opt}}}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}\log(q+1)}{2(q+1)^{1-w}}\left(1+16\sqrt{\log\left(\frac{8(q+2)^{2}}{\delta}\right)}\right)t^{-w}$
$\quad+2\varsigma\sup_{z\in X}|F(z)|t^{-p}+\left(\varsigma(q+1)^{\omega-p}+\varsigma\frac{(t+1)^{1-p}-q^{1-p}}{1-p}\right)t^{-w}\sup_{z\in X}|\mathcal{F}(z)|,$
where $\beta_{i}$’s are defined in (4) and $a,b$ are constants that satisfy
$\varsigma\frac{(q+2)^{w-p}}{3}+\varsigma\frac{t^{1-p}-(q+1)^{1-p}}{1-p}=a+bt^{1-p}.$
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}\!=\!\alpha$ for any $0\leq t\leq q$ and $\alpha_{t}=(t+1)^{-\omega}$ with $\omega>\frac{1}{2}$ for any $t>q$. If ˜1 holds with the exception that $F$ may be nonconvex, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}$ (28)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)},$
@@
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(8/\delta\right)}$ (29)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t^{1-2w}-s_{t}^{1-2w})}{S(1-2w)}}\sqrt{\log\left(8/\delta\right)},$
@@
Proof of Lemma˜14.
Recall from the proof of Lemma˜9 that, employing the concentration inequality in [41, Theorem 3.5] yields
@@
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $\delta/2$ and solving for $\lambda$ yields
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}$
@@
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\sqrt{\frac{\alpha^{2}q}{S}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}.$
@@
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 over the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha$ for any $0\leq t\leq q$ and $S=q$. If ˜1 holds with the exception that $F$ may be nonconvex and ˜4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, for any $0\leq t\leq q+1$, it holds that
@@
Proof of Lemma˜15.
Applying the bound (13) and after a straightforward re-arrangement, under ˜4, for any $0\leq i\leq t$, we obtain
@@
Given $\delta\in(0,1)$, we apply Lemma˜14 to deduce that with probability at least $1-\delta/(i+1)(i+2)$, we have
@@
We finally analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 after the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha$ for any $0\leq t<q$, $\alpha_{t}=(t+1)^{-w}$ with $\omega>\frac{1}{2}$ for any $t\geq q$ and $S=q$. If ˜1 holds with the exception that $F$ may be nonconvex and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}\frac{1}{(i+1)(i+2)}$, for any $t>q$, it holds that
@@
Given $\delta\in(0,1)$ and $i>q$, by Lemma˜14 and the inequality Eq.˜31, with probability at least $1-\delta/(i+1)(i+2)$, we have
@@
We now instantiate the regularization parameter and analyze Algorithm˜2.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\alpha:=\log(q+1)/(q+1)$ for any $0\leq t\leq q$, $\alpha_{t}=1/(t+1)^{\omega}$ for any $t>q$ and $S=q$. If ˜1 holds with the exception that $F$ may be nonconvex and ˜4 is satisfied, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<t-1}1/(i+1)(i+2)$, for any $t>q+1$, it holds that
@@
Proof of Lemma˜16.
By Lemma˜3, for any $t\geq q+1$, we have
@@
Applying Lemma˜15, with probability at least $1-\delta\sum_{0\leq i\leq q}1/(i+1)(i+2)$, we have
@@
Combining these two bounds with Proposition˜4 then yields
@@
We are now well-equipped to prove Theorem˜4.
Proof of Theorem˜4.
We first focus on the terms involving $\mathcal{F}(x_{i})$ in Lemma˜16. Note that the function $\mathcal{F}$ is defined with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
$-\sigma_{0}(q+1)^{\omega}\sum_{i=0}^{q}\alpha\mathcal{F}(x_{i})(1-\alpha)^{q-i}-\sum_{i=q+1}^{t-1}\sigma_{i}\mathcal{F}(x_{i})$
@@
where $\alpha=\log(q+1)/(q+1)$, and the second and third inequalities follow from the geometric series bound and the definition of $\sigma_{t}$. Using this bound together with Lemma˜16, we arrive at
$\mathcal{G}(x_{t})$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)t^{-w}$ (32)
@@
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Lemma˜16 and the definition of the sequence $\{\beta_{t}\}_{t\geq 0}$, we obtain
@@
$\frac{\sum_{i=0}^{t}\beta_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\beta_{i}}$ $\leq\left(\frac{\mathcal{G}(x_{0})}{(q+1)^{1-\omega}}+\left(2\varsigma(q+1)^{\omega-p}+\varsigma(q+2)^{-p}\right)\sup_{z\in X}|F(z)|\right)\left(a+bt^{1-p}\right)^{-1}$ (33)